OpenStudy (anonymous):

find the equation of a parabola with vertex (0,0) x- axis passing through point (-1,8)

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

y=a(x-h)^2+k 8=a(-1+0)^2+0 a=8 equation y=8x^2

OpenStudy (anonymous):

i did not see the answer in choose what I have

OpenStudy (anonymous):

this is what I have a) x^2=(1/8)y B) x^2=-8y c) (y-1)^2=4x D) y^2=16x

OpenStudy (anonymous):

a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u r wlcm